OpenStudy (anonymous):

An exective committee consists of 13 members: 6 men and 7women. 3 members are selected at random to attend an important meeting. The names are drawn from a hat. what is the probabilty that all 3 selected will be women?

hero (hero):

about 12.2% chance

hero (hero):

\[\frac{7}{13} \times \frac{6}{12} \times \frac{5}{11}\]